OpenStudy (anonymous):

what is the midpoint of the line segment whose endpoints are (-8,12) and (-13, -2)?

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -8}}\quad ,&{\color{blue}{ 12}})\quad &({\color{red}{ -13}}\quad ,&{\color{blue}{ -2}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)

OpenStudy (anonymous):

same as before take the average in each coordinate i.e. add and divide by two

OpenStudy (anonymous):

the answers I have are a.(-10.5, 5) b. ((-10.5, 7) c.(-2.5, 7)

OpenStudy (jdoe0001):

so.. what did you get then?

OpenStudy (anonymous):

c I think Im not good at this

OpenStudy (jdoe0001):

\(\bf \bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -8}}\quad ,&{\color{blue}{ 12}})\quad &({\color{red}{ -13}}\quad ,&{\color{blue}{ -2}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right) \\ \quad \\ \left(\cfrac{{\color{red}{ -13}} + {\color{red}{ -8}}}{2}\quad ,\quad \cfrac{{\color{blue}{ -2}} + {\color{blue}{ 12}}}{2} \right)\implies \left(\cfrac{-13-8}{2}\quad ,\quad \cfrac{-2+12}{2}\right)\) what do you think?

OpenStudy (anonymous):

B

OpenStudy (jdoe0001):

hmmm so... what would you get from the fractions?

OpenStudy (jdoe0001):

what's say -13 -18 = ? what about -2 + 12 = ?